OpenStudy (anonymous):

What is the equation of the line meeting the given conditions? Please put the equation in standard form. Containing A(5, 3) and perpendicular to a line with slope of -2

OpenStudy (anonymous):

slope of the line =1/2 y=(1/2)x+c putting (5,3) c=1/2 2y=x+1

OpenStudy (anonymous):

slope of th eperp line=-1/2 eqn is y-3=-1/2(x-5)

OpenStudy (anonymous):

2y=-x+11

OpenStudy (anonymous):

Thanks!